Find the critical value for c = 0.95 and n = 10Find the critical value t, for c = 0.95 and n = 10. 02.262 0 2.228 01.960 02.201
Answer #1

solution

Degrees of freedom = df = n - 1 = 10- 1 = 9

At 0.95 confidence level the t is ,

\alpha =1 - 0.95 = 0.05

  \alpha/ 2= 0.05 / 2 = 0.025

t\alpha /2,df = t0.025,9 =2.262 ( using student t table)